OpenStudy (anonymous):

Hello! Could somebody tell me how can I linearizing a second order non-linear equation? So I have f(x,y)= xy. Can I make a linear approximation on the f(x,y) function? Thank you!

OpenStudy (anonymous):

y = f(a) + f '(a) (x - a)

OpenStudy (anonymous):

is this tangent to f ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you!

hartnn (hartnn):

its actually, \(\large f\left(x,y\right)\approx f\left(a,b\right)+\frac{\partial f}{\partial x}\left(a,b\right)\left(x-a\right)+\frac{\partial f}{\partial y}\left(a,b\right)\left(y-b\right).\) for 2 variables.

hartnn (hartnn):

y = f(a) + f '(a) (x - a) applies only for f(x) as the function of x only....

hartnn (hartnn):

here u get \(\huge \frac{\partial f}{\partial x}=y \\\huge \frac{\partial f}{\partial y}=x \)

hartnn (hartnn):

so, \(\huge \frac{\partial f(a,b)}{\partial x}=b \\\huge \frac{\partial f(a,b)}{\partial y}=a \\ f(a,b)=ab\) now just plug in values.

hartnn (hartnn):

@silvia

OpenStudy (anonymous):

@hartnn thank you!

hartnn (hartnn):

welcome ^_^